OpenStudy (anonymous):

the absolute value of 5x-5/9 = 5

OpenStudy (anonymous):

-5=5x-5/9=5 -5+5/9=5x=5+5/9 -40/9=5x=50/9 -8/9=x=10/9

OpenStudy (anonymous):

thank you!